OpenStudy (anonymous):

Please help! thanku (-x^6)3 A.) -x^18 B.) -x^9 C.) x^3 D.) x^18

OpenStudy (anonymous):

A.) -x^18

OpenStudy (anonymous):

how did u do that?

OpenStudy (anonymous):

u multiply exponents

sam (.sam.):

@nick123411 Please explain, thank you. :)

OpenStudy (jdoe0001):

\(\large\bf (a^n)^m\implies a^{n\cdot m}\)

OpenStudy (anonymous):

(-x^6)3= -x^6*-x^6*-x^6= =-x^18

OpenStudy (anonymous):

(-x^6)3= -x^6*-x^6*-x^6= -x*x*x*x*x*x times -x*x*x*x*x*x times -x*x*x*x*x*x = -x*x*x*x*x*x* x*x*x*x*x*x *x*x*x*x*x*x = =-x^18